Rylee88:

How do I evaluate math? My equation is 3*4-9+12/2 12/2 is 12 divided by 2 NOT a fraction

Rylee88:

It's NOT 7.5

snowflake0531:

\(3\times4-9+\frac{12}{2}\) hint: think PEMDAS, first do all the multiplication and division things

Rylee88:

3*4 = 12

Rylee88:

12 divided by 2 = 6

snowflake0531:

so you have 12-9+6

Rylee88:

12-9=3+6=9

snowflake0531:

good job

Rylee88:

ahh okay ty.

snowflake0531:

yw~